In a store, out of all the people who came there 33 percent bought a shirt. If four people came in the store together and acted independently, find the probability that exactly one of them buys a shirt.

Round your answer to three decimal places.
Answer #1

Given that ,

p = 0.33

1 - p = 1 - 0.33= 0.67

n =4

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X =1 ) = (4 C 1) * 0.331 * (0.67)3

probability =0.3970